# Advancing Mobile GUI Agents: A Verifier-Driven Approach to Practical Deployment

### v4

@@ Abstract.
V-Droid obtains a substantial task success rate across several public mobile task automation benchmarks: 59.5% on AndroidWorld, 38.3% on AndroidLab, and 49% on MobileAgentBench, surpassing existing agents by 5.2%, 2.1%, and 9%, respectively. Furthermore, V-Droid achieves a remarkably low latency of 4.3s per step, which is $6.1\times$ faster compared with existing mobile agents. The source code is available at https://github.com/V-Droid-Agent/V-Droid.
The Work is done during internship at Microsoft Research
copyright: acmcopyrightcopyright: rightsretainedconference: Preprint;
@@ 1. Introduction
In recent years, a variety of LLM-powered mobile GUI agents have been proposed (Wen et al., 2024a; Agashe et al., 2025; Qin et al., 2025; Gou et al., 2025; Zheng et al., 2024; Wang et al., 2024c, e; Hong et al., 2024; Yang et al., 2024). These agents typically utilize LLMs as generators, generating decisions (e.g., reasoning and actions) based on the current task states (e.g., user interfaces, task descriptions), leveraging the contextual understanding and reasoning abilities inherent to LLMs. These agents, however, fall short of meeting practical deployments. Fig.1 presents the performance of state-of-the-art (SOTA) mobile agents in terms of task success rate (SR) and stepwise latency on the AndroidWorld benchmark(Rawles et al., 2024). As shown, the highest SR achieved by existing agents is only 54.3%, significantly lower than the human performance of 80%. Additionally, latency remains a critical challenge. Agent-S2 (Agashe et al., 2025), powered by GPT-4o and UI-TARS-72B (Qin et al., 2025), takes over 25 seconds for a single step.
@@ 4.1. Decision-Making Training
It is worth noting that, to the best of our knowledge, no publicly available dataset contains the required fine-grained process labels with contrastive action pairs. Therefore, we propose a novel data collection and synthesis approach, which will be detailed in Section 5.
@@ 4.2. Self-Correcting Training
It is worth noting that the number of erroneous states exceeds the number of correct states for a given task. We find training with excessive self-correcting pairs would lead to collapsed performance by continuing output navigating back. Therefore, we randomly sample the erroneous states to construct the self-correcting training pairs, which account for approximately 2.5% of the entire training set.
@@ 5.3. Training Details
Figure 9. Task success rate achieved by V-Droid compared to a range of mobile agents on three public benchmarks.
(a) Step-wise latency
(b) Decision-making latency
@@ 6.1.2. Baselines
V-Droid is compared with a wide range of mainstream mobile agents, including text-only agents (T3A (Rawles et al., 2024), AutoDroid (Wen et al., 2024a), AutoDroid-V2 (Wen et al., 2024b), MobileGPT (Lee et al., 2024), Ponder&Press (Wang et al., 2024e)), multimodal agents (M3A (Rawles et al., 2024), SeeAct (Zheng et al., 2024), AndroidArena (Xing et al., 2024), CogAgent (Hong et al., 2024), APPAgent (Li et al., 2024b), MobileAgent (Wang et al., 2024d)), and agents with grounding models (Agent-S2 (Agashe et al., 2025), Aria-UI (Yang et al., 2024), UGround (Gou et al., 2025), and UI-TARS (Qin et al., 2025)). Advanced LLMs, including GPT-3.5, GPT-4, GPT-4V, GPT-4o, GLM, DeepSeek-R1, Qwen, Llama-3.1, Claude and Gemini, are included as the base models for these agents.
@@ 6.2. Performance Improvement
Improved task success rate. Fig. 9 demonstrates that V-Droid outperforms existing mobile agents across three realistic MobileAgentBenchmarks, AndroidWorld, AndroidLab, and MobileAgentBench, achieving success rate (SR) improvements of 5.2%, 2.1%, and 9.0%, respectively. Compared to cloud-based LLM-powered agents (e.g., GPT-4, GPT-4o, DeepSeek-R1, Gemini-1.5-Pro, Claude-3.5), V-Droid achieves 25.0% and 7.13% higher SR on AndroidWorld and AndroidLab, respectively. Against advanced mobile agent frameworks (e.g., Agent-S2, AutoDroid, AppAgent), V-Droid improves SR by 5.2% on AndroidWorld and 9.0% on the MobileAgentBench. Compared to models that decompose decision-making into reasoning and grounding (e.g., UI-TARS, Aria-UI, UGround, Aguvis), V-Droid demonstrates a 14.3% SR improvement on the AndroidWorld. Against other fine-tuned SLMs (e.g., Qwen-VL-7B-FT, Llama-3.1-8B-FT), V-Droid achieves a notable 14.3% improvement on AndroidLab. Compared with memory-driven agents like AutoDroid and MobileGPT, V-Droid achieves over 36.2%, 21.0%, and 18.0% SR improvement on the three benchmarks.
@@
Figure 11. A showcase of V-Droid automating the task ’Check the file property of Textfile.txt’ on MobileAgentBench. Note that both this task and application were excluded from the training process
Comparison with CoT-Disabled Agents. The advantages of the verification-driven workflow in V-Droid are further demonstrated in Table 2 on AndroidWorld. Disabling CoT reasoning in GPT-4o and UI-TARS leads to reduced decision-making latency. However, this comes at the cost of decreased SR. Despite removing CoT, both GPT-4o and UI-TARS still exhibit higher latency than V-Droid, as they autoregressively generate full actions. In contrast, V-Droid achieves a substantially higher SR with lower decision-making latency. These results further underscore the generation– verification gap in GUI agents and highlight the efficiency and effectiveness of verifier-driven decision-making.
@@ 6.6. Failure Study
We conducted a manual analysis of the failure cases for V-Droid and MobileGPT across more than 300 tasks spanning three benchmarks. We identify four categories of failure: 1) Hallucinated decisions, where the agent selects an incorrect action despite being provided with the correct context; 2) Inaccurate memory, where erroneous contextual memory leads the agent to make wrong decisions; 3) Incomplete actions, wherein available action types are insufficient to complete the task; and 4) Modality limitations, where tasks cannot be accomplished by agents based solely on text-only LLMs.
As illustrated in Fig. 14, failure cases of MobileGPT primarily stem from suboptimal decision-making. For example, given the task instruction Add one expense (\$307.01, in Health Care categories) to APP Expense Pro, MobileGPT erroneously inputs the wrong amount and neglects to select the Health Care category, whereas V-Droid accurately provides all required information. This observation further underscores the necessity of post-training for GUI agents for a larger scale. Additional sources of failure include an incomplete action space caused by inaccurate information from the accessibility service, the lack of visual processing required for interpreting images and videos.
@@ 7. Discussion
Figure 14. Failure Analysis on MobileGPT and V-Droid.
Multimodality. Although V-Droid is currently text-only, the proposed approach can also be extended to train multimodal mobile agents. For instance, a grounding model could be trained to generate the initial actions, followed by the action verification and completion process.

### v5

# V-Droid: Advancing Mobile GUI Agent Through Generative Verifiers

@@ Abstract.
V-Droid obtains a substantial task success rate across several public mobile task automation benchmarks: 59.5% on AndroidWorld, 38.3% on AndroidLab, and 49% on MobileAgentBench, surpassing existing agents by 5.2%, 2.1%, and 9%, respectively. Furthermore, V-Droid achieves a remarkably low latency of 4.3s per step and 0.7s per decision, which are $6.1\times$ and $32.1\times$ faster compared with existing mobile agents. The source code is available at https://github.com/V-Droid-Agent/V-Droid.
Large Language Models, GUI Agent, Generative Verifier
journalyear: 2026copyright: ccconference: The 32nd Annual International Conference on Mobile Computing and Networking; October 26–30, 2026; Austin, TX, USAbooktitle: The 32nd Annual International Conference on Mobile Computing and Networking (MobiCom ’26), October 26–30, 2026, Austin, TX, USAdoi: 10.1145/3795866.3796681isbn: 979-8-4007-2505-0/2026/10ccs: Human-centered computing Ubiquitous and mobile computingccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
In recent years, a variety of LLM-powered mobile GUI agents have been proposed (Wen et al., 2024a; Agashe et al., 2025; Qin et al., 2025; Gou et al., 2025; Zheng et al., 2024; Wang et al., 2024c, e; Hong et al., 2024; Yang et al., 2024). These agents typically utilize LLMs as generators, generating decisions (e.g., reasoning and actions) based on the current task states (e.g., user interfaces, task descriptions), leveraging the contextual understanding and reasoning abilities inherent to LLMs. These agents, however, fall short of meeting practical deployments. Fig.2 presents the performance of state-of-the-art (SOTA) mobile agents in terms of task success rate (SR) and stepwise latency on the AndroidWorld (Rawles et al., 2024). As shown, the highest SR achieved by existing agents is only 54.3%, significantly lower than the human performance of 80%. Additionally, latency remains a critical challenge. Agent-S2 (Agashe et al., 2025), powered by GPT-4o and UI-TARS-72B (Qin et al., 2025), takes over 25 seconds for a single step.
@@ 4.1. Decision-Making Training
It is worth noting that, to the best of our knowledge, no publicly available dataset contains the required fine-grained process labels with contrastive action pairs. Therefore, we propose a novel data collection and synthesis approach, which will be detailed in §5.
@@ 4.2. Self-Correcting Training
It is worth noting that the number of erroneous states exceeds the number of correct states for a given task. We find training with excessive self-correcting pairs would lead to collapsed performance by continuing output the navigating back action. Therefore, we randomly sample the erroneous states to construct the self-correcting training pairs, which account for approximately 2.5% of the entire training set.
@@ 5.3. Training Details
Figure 9. Task success rate achieved by V-Droid compared to a range of mobile agents on three public benchmarks. The leftmost bar corresponds to V-Droid.
@@ 6.1.2. Baselines
V-Droid is compared with a wide range of mainstream mobile agents, including text-only agents (T3A (Rawles et al., 2024), AutoDroid (Wen et al., 2024a), AutoDroid-V2 (Wen et al., 2024b), MobileGPT (Lee et al., 2024), Ponder&Press (Wang et al., 2024e)), multimodal agents (M3A (Rawles et al., 2024), SeeAct (Zheng et al., 2024), AndroidArena (Xing et al., 2024), CogAgent (Hong et al., 2024), APPAgent (Li et al., 2024b), MobileAgent (Wang et al., 2024d)), and agents with grounding models (Agent-S2 (Agashe et al., 2025), Aria-UI (Yang et al., 2024), UGround (Gou et al., 2025), and UI-TARS (Qin et al., 2025)). Advanced LLMs, including GPT-3.5, GPT-4, GPT-4V, GPT-4o, GLM, DeepSeek-R1, Qwen, Llama-3.1, Claude and Gemini, are included as the base models for these agents. All the cloud-source LLMs are prompted using the M3A template in AndroidWorld (Rawles et al., 2024). Note that the baseline differences across benchmarks are attributable to the fact that some agents (Liu et al., 2024; Qin et al., 2025; Li et al., 2024b), were not open-sourced or lacked reproducible implementations, making it infeasible to evaluate them across all benchmarks. Therefore, we followed the practice and used all baselines available in each benchmark’s leaderboard to ensure fairness and completeness.
@@ 6.1.4. Evaluation platform
(a) Step-wise latency
(b) Decision-making latency
@@ 6.2. Performance Improvement
Improved task success rate. Fig. 9 demonstrates that V-Droid outperforms existing mobile agents across three realistic MobileAgentBenchmarks, AndroidWorld, AndroidLab, and MobileAgentBench, achieving success rate (SR) improvements of 5.2%, 2.1%, and 9.0%, respectively. Compared to cloud-based LLM-powered agents (e.g., GPT-4, GPT-4o, DeepSeek-R1, Gemini-1.5-Pro, Claude-3.5), V-Droid achieves 25.0% and 7.13% higher SR on AndroidWorld and AndroidLab, respectively. Against advanced mobile agent frameworks (e.g., Agent-S2, AutoDroid, AppAgent), V-Droid improves SR by 5.2% on AndroidWorld and 9.0% on the MobileAgentBench. Compared to models that decompose decision-making into reasoning and grounding (e.g., UI-TARS, Aria-UI, UGround, Aguvis), V-Droid demonstrates a 14.3% SR improvement on the AndroidWorld. Against other fine-tuned SLMs (e.g., Qwen-VL-7B-FT, Llama-3.1-8B-FT), V-Droid achieves a notable 14.3% improvement on AndroidLab. Compared with memory-driven agents like AutoDroid and MobileGPT, V-Droid achieves over 36.2%, 21.0%, and 18.0% SR improvement on the three benchmarks. Note that on MobileAgentBench, baselines such as AppAgent (Li et al., 2024b) already outperform cloud LLMs so those results are not included.
@@
Comparison with CoT-Disabled Agents. The advantages of the verification-driven workflow in V-Droid are further demonstrated in Table 2 on AndroidWorld. We compare V-Droid with the best agents using cloud-based GPT-4o and open-sourced local-served UI-TARS-7B (Qin et al., 2025) (similar size with V-Droid) that support execution with or without CoT. Other agents (e.g., Agent-S2 (Agashe et al., 2025)) rely on mandatory CoT or perform even worse, thus are excluded for this comparison. Disabling CoT reasoning in GPT-4o and UI-TARS leads to reduced decision-making latency. However, this comes at the cost of decreased SR. Despite removing CoT, both GPT-4o and UI-TARS still exhibit higher latency than V-Droid, as they autoregressively generate full actions. In contrast, V-Droid achieves a substantially higher SR with lower decision-making latency. V-Droid requires far fewer input tokens because it operates on simplified XML, lightweight system instructions, and compressed memory, whereas GPT-4o and UI-TARS depend on screenshots and excessive historical context. Besides, GPT-4o’s latency does not scale proportionally with output token reductions, which might due to the internet latency dominate its runtime. These results further underscore the generation–verification gap in GUI agents and highlight the effectiveness of verifier-driven workflow.
@@
Figure 11. A showcase of V-Droid automating the task ”Check the file property of Textfile.txt” on MobileAgentBench. Note that both this task and application were excluded from the training process
@@ 6.6. Failure Study
We conducted a manual analysis of the failure cases for V-Droid and MobileGPT across more than 300 tasks spanning three benchmarks. We identify four categories of failure: 1) Hallucinated decisions, where erroneous actions are taken despite the agent possessing accurate memory and full observational input; 2) Inaccurate memory, where erroneous contextual memory leads the agent to make wrong decisions; 3) Incomplete actions, wherein available action types are insufficient to complete the task; and 4) Modality limitations, where tasks that demands vision capabilities cannot be accomplished by text-only LLM agents. MobileAgentBench tasks inherently require minimal perception and memory, leading to a high percentage of failures in the hallucinated decision category.
As illustrated in Fig. 14, failure cases of MobileGPT primarily stem from suboptimal decision-making. For example, given the task instruction Add one expense (\$307.01, in Health Care categories) to APP Expense Pro, MobileGPT erroneously inputs the wrong amount and neglects to select the Health Care category, whereas V-Droid accurately provides all required information. There are still a large portion of failure cases of V-Droid in hallucinated decisions. For example, when required to share a file, V-Droid presses the file to read it instead of long-pressing it to reveal the sharing button, which might due to the lack of functional understanding of some UI elements. This observation further underscores the necessity of a larger scale training on more diverse tasks. Additional sources of failure of V-Droid include an incomplete action space caused by inaccurate information from the accessibility service and the lack of visual processing required for interpreting images and videos, which are further discussed in § 7.
@@ 7. Discussion
Figure 14. Failure Analysis of MobileGPT and V-Droid.
Multimodality. Although V-Droid is currently text-only, the proposed approach can also be extended to train multimodal mobile agents. For instance, V-Droid can be integrated with a grounding model with vision capability, which generate the initial actions for V-Droid instead of relying on accessibility services, followed by the action verification and completion process. We also plan to further train a vision-language model as the verifier as the future work, which assigns scores to actions based on the vision and text information jointly.
@@ 8. Conclusion
Acknowledgements.
This research is partially supported by Singapore Ministry of Education under its AcRF Tier 1 grant RT14/22, the Global STEM Professorship Scheme of Hong Kong, the HKUST start up grant, and the Research Grants Council (RGC) General Research Fund (GRF) 16210425.